Krissy525:

Using logs to solve exponential equations 2^x=10

Tbone:

Is it 2^x or x²

Krissy525:

2^x

Krissy525:

\[2^{x}\]

Tbone:

Okay, well have you tried this website called demos? If you just go you can put 2^2= 2^3= 2^4= 2^5= 2^6= until you get 10

Krissy525:

2^3 is 8 and 2^4 is 16

Tbone:

So that would make 2^5 10 right?

Krissy525:

no 2*5 is 10 2^5 is 32

Tbone:

2^2.5?

Krissy525:

5.657

Tbone:

Okay hang on lemme open up desmos

RyleyRoo22:

Step 1 of 2 : Write down the given equation. The given equation is. 2x = 10. Step 2 of 2 : Find the value of x. Dividing both sides by 2 we get. Hence the required solution is x = 5.

RyleyRoo22:

I believe this is right

Krissy525:

2 is raised to the power of x

Krissy525:

and 2 to the 5th is 32

RyleyRoo22:

Idk then Im a 12th grader so I dont use my brain for math

Krissy525:

lol thnx for trying tho

RyleyRoo22:

Your welcome